# Leveraging data-driven weather models for improving numerical weather prediction skill through large-scale spectral nudging

### v2

@@ 6.6 Computational cost of spectral nudging
The current implementation of spectral nudging in GEM with the presented optimal configuration leads to an increase of computational cost by approximately 25%. Applying nudging to specific humidity increases the computational cost by another 10% approximately. Due to the proof-of-concept nature of this study, no optimization of the informatics code has been performed. Therefore, the cost could be considerably reduced through future code optimization for the spectral filter and implementation of an asynchronous input server. Further improvements could be attained by considering time-varying nudging increments, where nudging would only be applied at time-steps close to the inference times of GC. These considerations however lie beyond the scope of this study.
@@ 7 Summary and Future Work
The recent emergence of AI models has begun to disrupt the operational weather forecasting paradigm which has long been dominated by the physics-based NWP models. Naturally, serious questions about the future role of AI and NWP models in operational forecasting are arising. For instance, are AI models capable of adding any reliable and meaningful value for weather services, or is it possible to improve guidance from operational NWP models by leveraging the AI models? This study attempts to address these pressing questions by choosing the GEM and GC models to represent the two forecasting approaches.
First, by comparing and contrasting the GEM and GC models, their relative strengths and weaknesses are identified. Skill verification against observations and GDPS analyses reveals that GC suffers from excessive fine-scale smoothing, affecting scales as large as 2500 km beyond day 3 of a 10-day forecast. However, despite smoothing, large-scale features (wavelength$>$2500 km) in GC exhibit better skill compared to GEM, which is evidenced by increased spectral coherence relative to the GDPS analyses and reduced STDE against radiosonde observations.
@@
Inspired by the improved large-scale skill of GC, a hybrid NWP-AI system – namely, GDPS-SN – has been developed to produce global forecasts where GEM-predicted large scales in horizontal wind and temperature are spectrally nudged toward inferences from GC. This hybrid system is capable of generating real-time forecasts with accuracy that surpasses ECCC’s operational GPDS. For example, RMSE of the 500-hPa geopotential height is reduced by 5-10%, with an overall predictability improvement of 12-18 hours, peaking at day 7 of a 10-day forecast. Notable improvements are also obtained in the predicted precipitation and tropical cyclone trajectories. Furthermore, predictions from GDPS-SN are expected to be more physically consistent than GC as, at its core, GDPS-SN employs an NWP model based on the first principles.
In addition to enhancing the skill and predictability of GEM, the proposed hybrid NWP-AI system is shown to generate the full spectra of fine scales that are missing from GC inferences. Comparisons against near-surface observations have revealed that – unlike GC, which has a limited ability to capture the extremes in wind speed and temperature – weather extremes predicted by GDPS-SN closely follow the operational GDPS-CTL.
Operational NWP models are generally required to track hundreds of prognostic variables while generating forecasts with high vertical resolution and temporal frequency. These capabilities are essential for meteorologists to forecast weather patterns, particularly those relevant for high-impact events. The proposed hybrid system maintains these necessary capabilities, albeit with a reasonable increase in computational cost. In this regard, it is important to recognize that similar capability requirements would substantially diminish the computational cost advantage of the AI models. Additionally, training such AI models would be prohibitively expensive.
Work is currently in progress at ECCC to fine-tune GC to emulate the operational GDPS analyses. This fine-tuned version is expected to enhance GC’s skill with the 37-level version, particularly in the stratosphere and the boundary layer. Any such improvement could be easily leveraged by the hybrid system in the near future. Importantly, a fine-tuned 37-level version of GC would allow for nudging in the boundary layer which is expected to result in further enhancement of near-surface skill with GDPS-SN, which with the current optimal configuration, shows less remarkable improvements over GDPS-CTL.
In the future, GDPS-SN could be utilized as the nudging reference for other downstream higher-resolution deterministic GEM-based operational NWP systems. Additionally, spectral nudging could be investigated for global ensemble prediction using a probabilistic AI model, such as GenCast (Price et al., 2024).
@@
Finally, results from this study imply that, rather than viewing AI and NWP models as two competing paradigms, it is likely more prudent to consider these forecasting methods as complementary. A well-designed fusion of these two approaches can significantly mitigate their individual limitations while allowing for harnessing their respective strengths to provide better meteorological guidance. Therefore, instead of focusing on efforts to replace one with the other, future research should prioritize improving both NWP and AI models.
@@ Acknowledgements.
where $n$ denotes the total wavenumber, $n_{o}$ is the cut-off total wavenumber and the exponent $r$ is a non-dimensional parameter related to the sharpness of the filter response. Based on the spectral comparison of the GDPS and GC predictions, the filter is configured by setting $n_{0}=30$ and $r=4$. As shown in Fig. 18, the resulting filtered fields fully retain amplitudes for scales associated with approximately $n<20$ (wavelength$>$2000 km) and fully removes scales corresponding to $n>40$ (wavelength$<$1000 km). The physical fields reconstructed from the filtered spectra leads to the desired filtered fields.
@@
Figures 19a and 19b illustrate the response of this DCT-based filter for the optimal configuration based on $\lambda_{LS}=2750$ km and $\lambda_{SS}=2250$ km. The figures demonstrate that the DCT-based filter is capable of targeting the desired scales for the individual Yin-Yang sub-domains. The response of the filter over the globe was determined by inspecting the spectral variance ratio of the filtered and unfiltered fields over a global Gaussian grid with respect to a spherical harmonics-based decomposition. The corresponding results are presented in Fig. 19c. The figure reveals that, even though DCT is primarily designed for LAMs where curvature effects of spherical surfaces are neglected, its application on global Yin-Yang grid does not result in any major issue as the filter is found to result in an acceptable response.
Figure 19: Response of the DCT-based filter (see Eq. 7) with respect to the (a) Yin grid, (b) Yang grid, and (c) a global Gaussian grid.

### v3

@@ 5.6 Computational cost of spectral nudging
The current implementation of spectral nudging in GEM with the presented optimal configuration leads to an increase of computational cost by approximately 25%. This is in addition to the relatively insignificant cost of generating the GraphCast inferences. The nudging overhead scales linearly with the number of fields and levels to which the technique is applied. Due to the proof-of-concept nature of this study, no optimization of the informatics code has been performed. In the future, this cost could be considerably reduced through code optimization for the spectral filter and implementation of an asynchronous input server. Further improvements could be attained by implementing time-varying nudging, where $\tau$ is allowed to change over time following a cosine-bell profile as proposed by Husain et al. (2014). With this approach, the application of spectral nudging could be restricted to model time steps close to the inference times of GraphCast, potentially leading to significant computational cost savings. However, these considerations lie beyond the scope of this study.
@@ 6 Summary and Future Work
The emergence of AI-based weather prediction models has disrupted the operational paradigm long dominated by the physics-based systems. The pressing question that motivates this study is whether it is possible to improve guidance from operational NWP models by leveraging the predictive skill of AI inferences.
@@
A careful comparison of physics-based GEM and AI-based GraphCast predictions reveals that the latter suffers from excessive smoothing up to synoptic scales. At larger scales, however, GraphCast predictions are found to be highly skillful. Inspired by the improved large-scale skill of GraphCast, a hybrid NWP-AI system – namely, GDPS-SN – has been developed to produce global forecasts in which GEM’s large-scale state is spectrally nudged toward larges scales of GraphCast inferences. This hybrid system is capable of generating real-time forecasts with accuracy that significantly surpasses ECCC’s operational GPDS. The RMSE of the 500-hPa geopotential height is reduced by 5-10%, with the largest predictability improvements attained around day 7 that exceed 24 hours over the tropics. This accurate prediction of large-scale circulation improves tropical cyclone steering estimates and the associated track predictions.
Although these results are achievable directly with AI-based systems, the hybrid model also generates the full spectra of fine scales that represents the tails of the climatology. This allows GDPS-SN to predict weather extremes that are challenging for the current generation of purely “data-driven” models.
Operational NWP models generate hundreds of internally and physically consistent forecast fields at high vertical resolution and temporal frequency. These outputs serve as essential guidance for operational meteorologists, especially for forecasts of high-impact events. The proposed hybrid system maintains these necessary capabilities, with a reasonable increase in computational cost. Increasing global AI model resolution and expanding the number of predicted variables represent real computational challenges for both training and inference. Moreover, the procedures needed to create data-driven predictions of the multitude of unanalyzed variables in a physically consistent way has not yet been devised. In this regard, the proposed hybrid system compensates for the perceived weaknesses of NWP models while addressing the limitations of current AI models.
The first version of GDPS-SN became operational (with an experimental status) at ECCC in March 2025. It is based on a GraphCast version that has been retrained and fine-tuned at ECCC. Work is also in progress to fine-tune GraphCast to emulate operational GDPS analyses. This fine-tuned version is expected to enhance GraphCast’s skill with the 37 pressure-level version, particularly in the stratosphere and the boundary layer. Any such improvement will quickly be integrated into the hybrid GDPS-SN system.
Although this study used the GEM and GraphCast models for hybridization, the protocol established here could be applied to any pair of systems, provided that the physical model supports spectral nudging. Hybrid systems like the one described here may represent an optimal blending of the individual advantages of the physics- and data-based approaches to weather prediction in the foreseeable future.
@@
The results from this study imply that, rather than viewing AI and NWP models as two competing paradigms, it is likely more prudent to consider these forecasting methods as complementary. A well-designed fusion of these two approaches can significantly mitigate their individual limitations while allowing for harnessing their respective strengths to provide better meteorological guidance. Therefore, instead of focusing on efforts to replace one with the other, future research should prioritize improving both NWP and AI models.
@@ Acknowledgements.
$A_{x}$ $=\sqrt{S_{x^{\prime},x^{\prime}}},$ (6)
$A_{y}$ $=\sqrt{S_{y^{\prime},y^{\prime}}},$ (7)
where, given any two 2-D scalar fields $u$ and $v$, the operator $S$ is defined as
$S_{u,v}\equiv E\bigg{[}\Big{\langle}{(u-\langle{u}\rangle)(v-\langle{v}\rangle)}\Big{\rangle}\bigg{]}.$ (8)
Here, the angle brackets denote spatial averaging, including latitudinal weighting.
The activity ratio of the forecast, normalized by the analysis activity, can then be obtained as follows:
$\Gamma=\sqrt{\frac{S_{x^{\prime},x^{\prime}}}{S_{y^{\prime},y^{\prime}}}},$ (9)
and the anomaly correlation coefficient is given as
$P=\frac{S_{x^{\prime},y^{\prime}}}{\sqrt{S_{x^{\prime},x^{\prime}}S_{y^{\prime},y^{\prime}}}}.$ (10)
These two coefficients quantify, respectively, the amplitude and phase errors of forecast anomalies with respect to the climatological mean.
By expressing the forecast and analysis fields as truncated spherical harmonic expansions, the covariance can be decomposed into contributions from individual harmonic components. Applying Parseval’s identity and grouping terms by zonal wavenumber, $S$ can be written as a sum over total spherical wavenumbers $n$ as follows:
$S_{u,v}=\sum_{n=1}^{N_{trunc}}E\big{[}\sigma_{u,v}(n)\big{]},$ (11)
where $\sigma_{u,v}(n)$ denotes the cross spectral density between $u$ and $v$ at wavenumber $n$. When $u=v$, $\sigma_{u,u}$ corresponds to the power spectral density of $u$. This allows for introducing a spectral activity ratio as a function of wavenumber $n$ as
$\gamma(n)=\sqrt{\frac{E\big{[}\sigma_{x^{\prime},x^{\prime}}(n)\big{]}}{E\big{[}\sigma_{y^{\prime},y^{\prime}}(n)\big{]}}},$ (12)
and a spectral anomaly correlation coefficient, hereafter referred to as spectral coherence, as
$\rho(n)=\frac{E\big{[}\sigma_{x^{\prime},y^{\prime}}(n)\big{]}}{\sqrt{E\big{[}\sigma_{x^{\prime},x^{\prime}}(n)\big{]}\,E\big{[}\sigma_{y^{\prime},y^{\prime}}(n)\big{]}}}.$ (13)
Estimating model climatology $x_{c}$ requires running hindcasts, which may be impractical outside of an operational NWP context. The conventional workaround is to use a common climatology database, setting $x_{c}=y_{c}=c$. The results presented in Figs. 19a–i correspond to this approach, where anomalies are calculated relative to a common ERA-Interim reanalysis climatology $c$ (Dee et al., 2011).
Alternatively, approximate climatological means $x_{c}$ and $y_{c}$ can be obtained as $x_{c}=\overline{x}$ and $y_{c}=\overline{y}$, where the overbar denotes average over a set of cases for a given season. This is the approach adopted throughout this paper. Therefore, Eqs. 1 and 2 are equivalent to Eqs. 12 and 13, respectively, where expectations are estimated by averaging spectral densities over a set of cases for a given season. Both approaches deliver qualitatively similar insights regarding model performance at different scales (not shown).
Figure 19: Normalized activity (left column), spectral activity ratio (middle column) at day 5, and scale-separated normalized activity ratio (right column) for temperature at 850 hPa (top row) and 250 hPa (middle row), and geopotential at 500 hPa (bottom row) for GDPS-CTL (blue), GDPS-SN (red), and GraphCast (grey). Forecast activity is calculated using ECMWF analysis climatology, whereas normalization is done using GDPS analyses. Results correspond to boreal winter.
Figure 19a illustrates the normalized activity ratio for temperature at 850 hPa across different lead times for predictions from GDPS-CTL, GDPS-SN, and GraphCast. By day 5, GraphCast’s activity is suppressed by about 4% compared to GDPS analyses. GDPS-CTL and GDPS-SN show similar activity, with GDPS-CTL being slightly overactive (up to 2%) beyond day 5. The spectral activity ratio at day 5 (Fig. 19b) shows increasingly suppressed activity for GraphCast at large wavenumbers, starting from wavenumber 16 (approx. 2500 km). For scales below 100 km, activity reduction is nearly 50%. However, this significant reduction in fine-scale activity is not reflected in the total activity ratio (Fig. 19a). Separating normalized activity for scales larger and smaller than 1000 km (wavenumber $<$40 and $>$40) in Fig. 19c shows that total activity ratio is mainly reflective of scales larger than 1000 km. This inference is based on the fact that the total activity ratios (dashed lines in Fig. 19c) almost match large-scale activity ratios (solid lines in Fig. 19a), while fine-scale activity ratios (dotted lines in Fig. 19c) indicate significant suppression in GraphCast.
Figures 19d–f further demonstrate the limitation of total activity ratio in identifying fine-scale smoothing effects. For temperature at 250 hPa, GraphCast’s total activity is the highest among models (Fig. 19d) despite a 60% reduction in activity for scales below 500 km (Fig. 19e). It turns out that the total activity metric rewards GraphCast for over-predicting activity at the planetary scales (wavenumber$<$4) and is insensitive to an overall 30% deficiency in activity for scales smaller than 1000 km (Fig. 19f).
Figures 19g–-i illustrate, for 500 hPa geopotential height, that activity as a metric can also be insensitive to spurious fine-scale variance. Specifically, both total activity (Fig. 19g) and fine-scale activity (Fig. 19i) for GraphCast fail to capture the unphysical increase in fine-scale variance concentrated near wavenumber 400 (Fig. 19h)—a known issue with GraphCast, as previously reported by Lam et al. (2023). Overall, these results highlight the limitations of total activity for assessing mesoscale issues in model predictions (smoothing or spurious variance) and underscores the importance of its spectral decomposition for proper assessment of model performance.
Finally, for all cases shown in Fig. 19, GDPS-SN’s fine-scale activity is close to GDPS-CTL, exhibiting only slight suppression depending on the specific variable and level. This behavior is consistent with the spectral activity and amplitude ratios shown in Figs. 19 and 7, respectively.
[…]
@@
where $n$ denotes the total wavenumber, $n_{o}$ is the cut-off total wavenumber, and the exponent $r$ is a non-dimensional parameter related to the sharpness of the filter response. Based on the spectral comparison of the GDPS and GraphCast predictions, the filter is configured by setting $n_{0}=30$ and $r=4$. As shown in Fig. 20, the resulting filtered fields fully retain amplitudes for scales associated with approximately $n<20$ (wavelength$>$2000 km) and fully removes scales corresponding to $n>40$ (wavelength$<$1000 km). The physical fields reconstructed from the filtered spectra leads to the desired filtered fields.
@@
Figures 21a and 21b illustrate the response of this DCT-based filter for the optimal configuration based on $\lambda_{LS}=2750$ km and $\lambda_{SS}=2250$ km. The figures demonstrate that the DCT-based filter is capable of targeting the desired scales for the individual Yin-Yang sub-domains. The global response of the filter was determined by first interpolating a field (e.g., temperature) from the Yin-Yang grid to a global Gaussian grid before and after applying the DCT-based filter. The spectral variance ratio between the filtered and unfiltered fields was then computed using a spherical harmonics-based decomposition. The corresponding results for temperature at 500 hPa are presented in Fig. 21c for illustration.
Figure 21: Response of the DCT-based filter based on $\lambda_{LS}=2750$ km and $\lambda_{SS}=2250$ km (see Eq. 15) with respect to the (a) Yin grid, (b) Yang grid, and (c) a global Gaussian grid for temperature at 500 hPa. Over the global Gaussian grid, the small- and large-scale cut-offs for the optimal configuration are approximately 2000 km and 5500 km, respectively, as indicated with the shaded area.